Five jobs are ready for processing at time zero through a workstation. The details of the processing time and the due date of completion is provided in the following table. Job Processing Time (days) Due Date A 6 18 B 3 14 C 4 9 D 5 19 E 2 11 Using the earliest due date (EDD) rule, the average flow time for the five jobs is _____. a. more than 10 but less than or equal to 12 days. b. more than 8 but less than or equal to 10 days. c. more than 12 but less than or equal to 14 days. d. more than 14 but less than or equal to 16 days.
Answer #1

According to the EDD rule the jobs are processed in order of their due date of completion. This means the job having the earliest due date is processed first and so on

In this problem job c has the earliest due date of 9 so it is proceesed first, the second is E with due date of 11,third is B with due date of 14,fourth is A with due date of 18 and last is D with due date of 19

So using the EDD rule the sequence of the job is C-E-B-A-D

Then we have to calculate the flow time of each job according to the sequence. Flow time = cumulative sum of the processing time

Job processing time flow time

C 4 4

E 2 6 (as 4+2 = 6)

B 3 9 (as 6+3 = 9)

A 6 15 (as 9+6 = 15)

D 5 20 (as 15+5 = 20)

Average flow time = sum of the flow time for each job/number of jobs

= (4+6+9+15+20) /5

= 54/5

= 10.8

So the answer is option a i. e more than 10 but less than or equal to 12 days